Algebra

If the remainders of the polynomial f(x) when divided by x – 1, x – 2 are 2, 5 then the remainder of f(x) when divided by (x – 1) (x – 2) is

rushanth
12 Years agoGrade

1 Answer

Sunil Raikwar
ApprovedApproved Tutor Answer12 Years ago
let the remainder be (ax+b) then
f(x)= (x-1)(x-2).q(x)+ax+b...................1
put x=1 & 2 in 1, we find
a+b=2
2a+b=5
a=3,b=-1
Required remainder is 3x-1

Thanks & Regards
Sunil Raikwar
askIITians Faculty